# Juntao Zhao1,*    Wenhao Lu2,†    Sheng Wang1,†    Lingpeng Kong1    Chuan Wu1

### v2

@@ Abstract
Quantization has been substantially adopted to accelerate inference and reduce memory consumption of large language models (LLMs). While activation-weight joint quantization speeds up the inference process through low-precision kernels, we demonstrate that it suffers severe performance degradation on multi-step reasoning tasks, rendering it ineffective. We propose a novel quantization paradigm called QSpec, which seamlessly integrates two complementary quantization schemes for speculative decoding. Leveraging nearly cost-free execution switching, QSpec drafts tokens with low-precision, fast activation-weight quantization, and verifies them with high-precision weight-only quantization, effectively combines the strengths of both quantization schemes. Compared to high-precision quantization methods, QSpec empirically boosts token generation throughput by up to 1.64×1.64$\times$1.64 × without any quality compromise, distinguishing it from other low-precision quantization approaches. This enhancement is also consistent across various serving tasks, model sizes, quantization methods, and batch sizes. Compared to state-of-art speculative decoding methods, our approach reuses weights and the KV cache, avoiding extra memory overhead while achieving up to 1.55×1.$5$ speedup in batched serving with a high acceptance rate. Furthermore, QSpec offers a plug-and-play advantage without requiring any training. We believe that QSpec demonstrates unique strengths for future deployment of high-fidelity quantization schemes, particularly in memory-constrained scenarios (e.g., edge devices).
@@ 1   Introduction
Large language models (LLMs) have demonstrated remarkable abilities across various domains, including mathematics, coding, and planning (Shao et al., 2024b; Guo et al., 2024a; Huang et al., 2024). Nonetheless, their immense scales pose substantial challenges for deployment due to high memory and computational demands, especially in resource-limited scenarios (e.g., inference on edge devices). Quantization has been an effective compression technique to facilitate LLM inference with limited resources (Lin et al., 2024a; Ashkboos et al., 2024; Zhao et al., 2024b; Lin et al., 2024b). By converting high-precision values (e.g., FP16) into their lower-precision counterparts (e.g., INT4), quantization effectively lowers memory and computational requirements, allowing for larger serving batches and model sizes. Furthermore, the reduced memory footprint boosts token generation throughput by accelerating the typically memory-bound autoregressive decoding process (Zhao et al., 2024a).
@@
“Is there a quantization solution that boosts efficiency while avoiding performance degradation?”.
@@
Figure 1: Diagrams of different 4-bit quantization schemes. Left: W4A16 uses 4-bit weight and 16-bit activation for inference. Middle: W4A4 further adopts 4-bit activation to utilize low-precision W4A4 kernels. Right: QSpec accelerates W4A16 by drafting tokens with W4A4 and verifying them with W4A16, and applies KV cache overwriting for consistent memory consumption.
Then to answer the above question, we draw inspiration from Speculative Decoding (Leviathan et al., 2023; Chen et al., 2023), which combines rapid drafting of a small model with high-quality token generation of a larger model to boost throughput (i.e., efficiency) without compromising performance (i.e., efficacy). We propose a novel paradigm called QSpec, which combines mixed-precision quantization execution to tackle the trade-off between efficiency and efficacy while maintaining the memory usage of high-precision quantization. Our key insight is that a single weight-quantized model can efficiently toggle two parallel quantization schemes: one with quantized activations and the other without, which we further empirically verify to produce highly similar tokens (Sec. 2.2). This observation unveils the potential for a synergistic approach combining both schemes. As illustrated in Figure 1, for a 4-bit weight-quantized model, we can leverage the faster yet lower-quality execution flow (i.e., W4A4) to draft tokens, while verifying these drafted tokens with the higher-quality quantization flow (i.e., W4A16) with negligible switching costs. Similar to speculative decoding, this ‘draft-verify’ mode with mixed quantization execution ensures high fidelity with the verifying flow. Differently, our approach re-utilizes the weights and high-precision KV cache, maintaining the memory overhead equivalent to that of the high-precision scheme alone, rather than the sum of both schemes in speculative decoding.
We evaluate the generation quality and end-to-end serving throughput of QSpec against W4A4 and W4A16 schemes across multiple datasets, model sizes, quantization methods, and batch sizes. Empirically, QSpec preserves memory consumption and generation quality compared to W4A16, while offering a high acceptance rate and up to 1.64×1.64$\times$1.64 × higher token generation throughput, thereby mitigating the efficiency-efficacy trade-off of existing quantization methods. Notably, for multi-step reasoning tasks such as MATH (Hendrycks et al., 2021), QSpec fully compensates for an up to 51.11% decline in generation quality observed with existing W4A4 methods. Furthermore, QSpec provides plug-and-play compatibility without any training requirements, and can be seamlessly applied to any existing models, delivering superior performance with minimal effort.
In summary, our main contributions are as follows:
  1. We demonstrate that multi-step reasoning tasks can better capture the performance variations of quantization schemes than current evaluation protocols, and advocate for their incorporation for more comprehensive assessment.
@@
  3. We propose QSpec, the first quantization paradigm that synergizing two complementary weight-shared quantization schemes with speculative decoding, alleviating the efficiency-efficacy trade-off of quantization. We demonstrate its superior efficiency in quantized batch-serving compared to state-of-the-art speculative decoding approaches theoretically and empirically.
  4. Our empirical results reveal up to 1.64×1.64$\times$1.64 × acceleration without any quality sacrifice across diverse settings. Alongside consistent memory usage, these advantages promise QSpec for high-fidelity quantization deployment, especially in memory-constrained scenarios.
@@ 2.1   Compromised Performance of Activation Quantization
State-of-the-art (SOTA) activation-weight joint quantization methods, like Atom (Zhao et al., 2024b) and QuaRot (Ashkboos et al., 2024), achieve notable speed-ups with negligible performance loss compared to weight-only ones. However, we argue that this conclusion is skewed by the limited evaluation benchmarks, which fail to capture the negative impacts of activation quantization.
To substantiate this claim, we conduct experiments on Llama-3-8B-Instruct models (Dubey et al., 2024) quantized with W16A16, W4A16, and W4A4 methods across four task datasets: PIQA (Bisk et al., 2020), WikiText-2 (Merity et al., 2016), GSM8K (Cobbe et al., 2021), and MBPP (Austin et al., 2021). PIQA is a two-choice commonsense reasoning benchmark for physical knowledge, evaluated using classification accuracy, while WikiText-2 comprises a collection of high-quality Wikipedia articles, assessed for language fluency via perplexity (Jelinek et al., 1977). Both are commonly adopted in current quantization evaluations. GSM8K includes diverse grade school mathematical problems, evaluated by “exact match” metrics; MBPP focuses on crowd-sourced Python programming challenges, assessed by accuracy. Unlike the former two benchmarks, both GSM8K and MBPP necessitate auto-regressive multi-step reasoning abilities. While these critical abilities are propelled by the rapid advancement in LLMs recently, they have not yet been widely integrated into mainstream quantization evaluation. As shown in Table 1, Atom-based quantization schemes show comparable performance to W16A16 across commonly adopted tasks such as on PIQA and WikiText-2, aligning with the claims in Zhao et al. (2024b). However, W4A4 suffers a nearly 30% average performance decline on complex reasoning tasks (i.e., on MBPP and GSM8K), whereas W4A16 only experiences about 4%. This indicates that activation quantization leads to several times more performance degradation on multi-step reasoning tasks, despite the improved efficiency. Besides, the performance trend observed on multi-step reasoning tasks shows a stronger correlation with quantization precision than perplexity does, validating their adequacy in quantization evaluation.
In summary, activation quantization still incurs significant performance loss on more advanced multi-step reasoning tasks. This necessitates the inclusion of reasoning tasks in quantization evaluation for a more comprehensive assessment. On the other hand, this also underscores the demand for a quality-preserving yet efficient quantization paradigm.
@@ 2.2   High-Similarity Token Predictions
Despite the notable performance decline caused by activation quantization, we observe, more microscopically, high similarity in top-1 token predictions between quantization schemes with high and low precision activations. Specifically, we first employ Atom-based W4A16 greedy sampling to generate the golden token sequences for the GSM8K test set, obtaining the prediction probabilities for each top-1 answer token. Subsequently, we perform one Atom-based W4A4 forward pass (i.e., prefill) on the concatenated input of each question and its corresponding golden answer to acquire the token probabilities as well. This allows us to assess the prediction discrepancy between W4A4 and W4A16. As illustrated in Figure 2, we observe that (1) the majority of token prediction probabilities of both W4A4 and W4A16 exceed 80%, and most of the tokens associated with high probabilities are accepted. (2) Compared to accepted tokens, the number of rejected ones is negligible, underscoring the high similarity between the two quantization methods. Combined with the analysis in Sec. 2.1, this can be interpreted that a small set of salient token variations can trigger a snowball effect of errors, especially on multi-step reasoning tasks where the subsequent steps are closely conditioned on the previous ones, akin to findings in Zhang et al. (2023), thus impairing the performance of the low-precision activation scheme. Prior studies indicate that low similarity leads to frequent token rejections, thereby diminishing the efficiency of speculative decoding (Leviathan et al., 2023). The observed high token-level similarity suggests that we could potentially restore the generation quality by detecting and correcting a limited number of generation errors incurred by activation quantization. This insight motivates us to propose a quantization-specific speculative decoding framework.
@@ 3   Method
Targeting an efficient quantization scheme without sacrificing performance or increasing memory consumption, we propose a new quantization paradigm called speculative decoding with complementary quantization execution (QSpec). As shown in Figure 1, QSpec employs a draft-verify pipeline for next-token prediction with varying activation precisions and shared low-precision quantized weights, instead of a single quantization scheme. The details are elaborated below, adhering to the core component breakdown of regular speculative decoding (Leviathan et al., 2023).
@@ 3.1   QSpec
Figure 3: A mini-sample of QSpec, where green, red, and purple tokens represent draft tokens, rejected tokens, and tokens generated directly by W4A16, respectively.
@@
Current LLMs typically utilize an autoregressive process for next-token prediction, where a new token is drawn from a probability distribution conditioned on all previously generated tokens. This process can be formulated as:
@@
Compared with previous research (Leviathan et al., 2023; Chen et al., 2023), on one hand, we employ a weight-shared quantization scheme with low-precision activations, rather than one standalone small-sized model, to speculate the next $\gamma$ tokens $\hat{T}_{i+1:i+\gamma}$ and their associated distributions $\hat{p}_{i+1:i+\gamma}(t)$. In $\hat{T}_{i+1:i+\gamma}$, each token $\hat{t}_{j}$ is sampled from $\mathcal{M}_{l}(\hat{t}_{j}|T_{\leq i},\hat{T}_{i+1:j-1})$, where $j\in[i+1,i+\gamma]$ and $\mathcal{M}_{l}$ represent our quantized model executed with low-precision activation. On the other hand, our low-precision quantization scheme shares similar attributes as the draft model in Leviathan et al. (2023), as both can generate tokens rapidly, though with reduced quality.
@@
To maintain high reproducibility, both low-precision and high-precision activation quantization schemes employ greedy decoding. This means that one drafted tokens $\hat{t}_{i+j}$ is accepted as ${t}_{i+j}$ only when the top-1 tokens from ${p}_{i+j}$ and $\hat{p}_{i+j}$ coincide; otherwise, this token is rejected. Nonetheless, we claim that alternative strategies, as outlined in Leviathan et al. (2023), can be directly applied to our method due to the similarities in the framework. Figure 3 illustrates a mini-sample of this cycle with the draft token length $\gamma=4$. The model initially speculates four tokens using W4A4 scheme. Subsequently, adhering to a predefined acceptance policy, it accepts all drafted tokens after verifying them through the W4A16 scheme. In the second loop, however, only the first two tokens are accepted. A new token “is” is directly derived from the prediction probability of W4A16 scheme, and another draft-verify cycle will commence from the ninth token.
@@
To further reduce memory consumption, QSpec overwrites KV caches of low-precision activation quantization with those of high-precision method. Specifically, compared to W4A4, W4A16 is expected to yield a higher quality FP16 KV cache due to higher activation precision. Alongside the shared weights, this naturally allows overwriting the low-quality KV caches generated by W4A4 with those from W4A16 for accepted tokens after each validation phase. This enables W4A4 to condition on high-quality KV caches for subsequent autoregressive generation, and saves the memory occupation of W4A4 KV caches, despite requiring negligible buffer space for temporarily storing. To some extent, this operation aligns with the settings of attention kernels in prior works (Zhao et al., 2024b; Shao et al., 2024a; Ashkboos et al., 2024), where INT4 KV caches are typically dequantized to FP16 before or during precision-sensitive attention operations to ensure accurate computations. Note that KV compression (Hooper et al., 2024; Kang et al., 2024) is fully orthogonal to our method, e.g., it is also possible to use 4-bit KV cache for W4A16.
Figure 4: Analysis of QSpec advantages: (a) Tree-structured drafting architecture, (b) Roofline model showing computational efficiency transitions (memory-bound $\rightarrow$ compute-bound) with batch size scaling, and (c) Per-valid-token latency composition comparing the change of extra computational overhead and accepted token numbers with scaling batch size. SD denotes tree-structured speculative decoding.
@@
QSpec
As shown in Table 2, we compare QSpec against individual quantization configurations (i.e., W4A4 and W4A16) and speculative decoding in terms of memory, computation, and generation. QSpec provides four key benefits:
(1) Memory-Efficient. By sharing weights and overwriting KV caches, QSpec incurs memory costs on par with standalone high-precision activation quantization. (2) No Efficiency–Efficacy Trade-off. QSpec leverages speculative decoding to boost efficiency while preserving output quality, avoiding the compromises seen in conventional quantization. (3) Plug-and-Play Compatibility. QSpec requires only an acceptance policy and a cache-overwriting step, requiring no additional training or classifiers. Hence, it can be quickly integrated into existing quantization pipelines. (4) High Acceptance Rate. Common weights and KV cache overwriting ensure consistent predictions, leading to a high rate of token acceptance.

### v3

# QSpec: Speculative Decoding with Complementary Quantization Schemes

@@ Abstract
Quantization is widely adopted to accelerate inference and reduce memory consumption in large language models (LLMs). While activation-weight joint quantization enables efficient low-precision decoding, it suffers from substantial performance degradation on multi-step reasoning tasks. We propose QSpec, a novel quantization paradigm that decouples efficiency from quality by integrating two complementary schemes via speculative decoding: low-precision joint quantization for fast drafting and high-precision weight-only quantization for accurate verification. QSpec reuses both weights and KV cache across stages, enabling near-zero-cost switching without retraining or auxiliary models. Compared to high-precision baselines, QSpec achieves up to $1.64\times$ speedup without quality degradation, and outperforms state-of-the-art speculative decoding methods by up to $1.55\times$ in batched settings. Furthermore, QSpec supports plug-and-play deployment and generalizes well across model scales, quantization methods, and workloads. These properties make QSpec a practical and scalable solution for high-fidelity quantized LLM serving under memory-constrained scenarios. Our code is available at https://github.com/hku-netexplo-lab/QSpec.
QSpec: Speculative Decoding with Complementary Quantization Schemes
Juntao Zhao, Wenhao Lu, Sheng Wang, Lingpeng Kong, and Chuan Wu {juntaozh, whlu, u3009618}@connect.hku.hk, {lpk, cwu}@cs.hku.hk
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable abilities across various domains, including mathematics, coding, and planning (Shao et al., 2024; Guo et al., 2024a; Huang et al., 2024). Nonetheless, their immense scales pose substantial challenges for deployment due to high memory and computational demands, especially in resource-limited scenarios (e.g., inference on edge devices). Quantization has been an effective compression technique to facilitate LLM inference with limited resources (Lin et al., 2024a; Ashkboos et al., 2024; Zhao et al., 2024b; Lin et al., 2024b). By converting high-precision values (e.g., FP16) into their lower-precision counterparts (e.g., INT4), quantization effectively lowers memory and computational requirements, allowing for larger serving batches and model sizes. Furthermore, the reduced memory footprint boosts token generation throughput by accelerating the typically memory-bound autoregressive decoding process (Zhao et al., 2024a).
Figure 1: Diagrams of different 4-bit quantization schemes. W4A16: uses 4-bit weight and 16-bit activation for inference. W4A4: further adopts 4-bit activation to utilize low-precision W4A4 kernels. QSpec: accelerates W4A16 by drafting tokens with W4A4 and verifying them with W4A16, and applies KV cache overwriting for consistent memory consumption.
@@
“Is there a quantization solution that boosts efficiency while avoiding efficacy degradation?”.
@@
To answer the above question, we draw inspiration from speculative decoding  (Leviathan et al., 2023; Chen et al., 2023), which combines rapid drafting of a small model with high-fidelity verification from a larger model to boost throughput without sacrificing output quality. We propose a novel paradigm called QSpec, which employs complementary mixed-precision quantization execution to deliver flawless acceleration for high precision quantization. (i.e., efficiency improvements that preserve the fidelity and memory overhead of high-precision quantization).
Our key insight is that a single weight-quantized model can losslessly toggle between two parallel activation modes: a fast, low-precision mode for drafting (e.g., W4A4) and a high-precision mode for verification (e.g., W4A16). As validated in Sec. 2.2, the token generation with different modes is highly similar. This allows us to adopt a lightweight ‘draft-verify’ scheme, as shown in Figure 1, where tokens drafted with W4A4 are selectively accepted by W4A16 with negligible switching costs. Unlike traditional speculative decoding that requires a draft model, QSpec shares weights and KV cache across activation modes, achieving zero extra memory overhead.
We evaluate QSpec on a range of model sizes, quantization methods, and batch sizes. Compared to W4A16, QSpec offers up to $1.64\times$ higher token generation throughput while preserving fidelity. It also effectively compensates for up to 51.11% quality loss observed in W4A4 on challenging multi-step reasoning tasks such as MATH (Hendrycks et al., 2021). Furthermore, QSpec requires no training and can be directly integrated into existing inference pipelines.
Our main contributions are summarized as follows:
  1. We demonstrate that multi-step reasoning tasks are more sensitive to quantization-induced quality degradation than standard benchmarks, and advocate their inclusion for more comprehensive evaluation.
@@
  3. We introduce QSpec, the first quantization paradigm that decouples efficiency from quality by combining complementary quantization schemes through speculative decoding with shared weights and KV cache.
  4. Experiments across various models and tasks reveal that QSpec achieves up to $1.64\times$ acceleration without quality loss, making it well-suited for high-fidelity deployment in memory-constrained scenarios.
@@ 2.1 Compromised Performance of Activation Quantization
State-of-the-art (SOTA) activation-weight joint quantization methods, like Atom (Zhao et al., 2024b) and QuaRot (Ashkboos et al., 2024), achieve notable speed-ups with negligible performance loss compared to weight-only ones. However, we argue that this conclusion is skewed by limited evaluation benchmarks, which fail to capture the negative impacts of activation quantization.
@@
To substantiate this claim, we conduct experiments on Llama-3-8B-Instruct models (Dubey et al., 2024) quantized with W16A16, W4A16, and W4A4 across four benchmarks: PIQA (Bisk et al., 2020), WikiText-2 (Merity et al., 2016), GSM8K (Cobbe et al., 2021), and MBPP (Austin et al., 2021). While PIQA and WikiText-2 are commonly used in quantization evaluation, GSM8K and MBPP involve multi-step reasoning, which remains underexplored in the quantization context despite its great importance. Detailed descriptions of the benchmarks are provided in Appendix B.
As listed in Table 1, Atom-based quantization schemes show comparable performance to W16A16 across commonly adopted tasks such as on PIQA and WikiText-2, aligning with the claims in Zhao et al. (2024b). However, W4A4 suffers a nearly 30% average performance decline on complex reasoning tasks (i.e., MBPP and GSM8K), whereas W4A16 only experiences about 4%. This indicates that activation quantization leads to several times more performance degradation on multi-step reasoning tasks, despite the improved efficiency. Besides, the performance trend observed on multi-step reasoning tasks shows a stronger correlation with quantization precision than perplexity does, validating their adequacy in evaluation.
In summary, activation quantization still incurs significant performance loss on more advanced multi-step reasoning tasks. This necessitates the inclusion of them in quantization evaluation for a more comprehensive assessment. Furthermore, this also underscores the demand for a quality-preserving yet efficient quantization paradigm.
@@ 2.2 High-Similarity Token Predictions
Despite the notable performance decline caused by activation quantization, we observe, more microscopically, high similarity in top-1 token predictions between quantization schemes with high and low precision activations. Specifically, we first employ Atom-based W4A16 greedy sampling to generate the golden token sequences for the GSM8K test set, obtaining the prediction probabilities for each top-1 answer token. Subsequently, we perform one Atom-based W4A4 forward pass (i.e., prefill) on the concatenated input of each question and its corresponding golden answer to acquire the token probabilities as well. This allows us to assess the prediction discrepancy between W4A4 and W4A16. As illustrated in Figure 2, we observe that (1) the majority of token prediction probabilities of both W4A4 and W4A16 exceed 80%, and most of the tokens associated with high probabilities are accepted. (2) Compared to accepted tokens, the number of rejected ones is negligible, underscoring the high similarity between the two quantization methods.
@@
Combined with the analysis in Sec. 2.1, this can be interpreted that a small set of salient token variations can trigger a snowball effect of errors, especially on multi-step reasoning tasks where the subsequent steps are closely conditioned on the previous ones, akin to findings in Zhang et al. (2023), thus impairing the performance of the low-precision activation scheme. Prior studies indicate that low similarity leads to frequent token rejections, thereby diminishing the efficiency of speculative decoding (Leviathan et al., 2023). The high token-level similarity we observe implies that generating high-quality outputs may only require detecting and correcting a limited number of activation quantization-induced errors. This insight motivates our proposal of a quantization-aware speculative decoding framework that leverages token generation similarity.
@@ 3 Method
Figure 3: A mini-sample of QSpec, where yellow, red, and blue tokens represent W4A4 draft tokens, rejected tokens, and tokens generated directly by W4A16, respectively. While these green ones are draft tokens that have been verified and accepted by W4A16.
Targeting an efficient quantization scheme without sacrificing performance or increasing memory consumption, we propose a new quantization paradigm called Speculative decoding with complementary Quantization execution (QSpec). As shown in Figure 1 and Figure 3, QSpec employs a draft-verify pipeline for next-token prediction with varying activation precisions and shared low-precision quantized weights, instead of a single quantization scheme. Only quantization schemes are switched, and no additional weights are incorporated in this process.
@@ 3.1 QSpec
Current LLMs typically utilize an autoregressive process for next-token prediction. A new token is drawn from a probability distribution conditioned on all previously generated tokens. This process can be formulated as:
@@
Compared with previous research (Leviathan et al., 2023; Chen et al., 2023), we employ a weight-shared quantization scheme with low-precision activations, rather than a standalone small-sized model, to speculate the next $\gamma$ tokens $\hat{T}_{i+1:i+\gamma}$ and their associated distributions $\hat{p}_{i+1:i+\gamma}(t)$. In $\hat{T}_{i+1:i+\gamma}$, each token $\hat{t}_{j}$ is sampled from $\mathcal{M}_{l}(\hat{t}_{j}|T_{\leq i},\hat{T}_{i+1:j-1})$, where $j\in[i+1,i+\gamma]$ and $\mathcal{M}_{l}$ represents our quantized model executed with low-precision activation. Thanks to the reduced activation precision, this scheme enables fast token generation.
@@
To maintain high reproducibility, both low-precision and high-precision activation quantization schemes employ greedy decoding. This means that one drafted tokens $\hat{t}_{i+j}$ is accepted as ${t}_{i+j}$ only when the top-1 tokens from ${p}_{i+j}$ and $\hat{p}_{i+j}$ coincide; otherwise, this token is rejected. Nonetheless, we claim that alternative strategies, as outlined in Leviathan et al. (2023), can be directly applied to our method due to the similarities in the framework. Figure 3 illustrates a mini-sample of this cycle with the draft token length $\gamma=4$. The model initially speculates four tokens using the W4A4 scheme. Subsequently, adhering to a predefined acceptance policy, it accepts all drafted tokens after verifying them through the W4A16 scheme. In the second loop, however, only the first two tokens are accepted. A new token “is” is directly derived from the prediction probability of the W4A16 scheme, and another draft-verify cycle will commence from the ninth token.
@@
A key advantage of QSpec lies in its shared-weight architecture, which naturally aligns the behavior of low- and high-precision activations. This allows the high-precision verification stage to produce activation patterns and KV cache values that serve as high-fidelity substitutes for those generated during the low-precision draft stage. Leveraging this alignment, QSpec overwrites the lower-quality KV caches from W4A4 with accurate A16 caches from W4A16 for accepted tokens, enabling subsequent decoding to benefit from higher-quality context. This design not only boosts token acceptance rates but also sets QSpec apart from prior speculative decoding methods, which use separate models and cannot reuse KV representations. By sharing weights and reusing KV caches within a single model, QSpec eliminates the need for dual cache maintenance, reducing memory usage without sacrificing accuracy.
@@
QSpec (no-overwrite) ✗ (1x) ✓ (1.25x) ✗ (0.8x)
QSpec ✗ (1x) ✗ (1x) ✓ (1x)
As shown in Table 2, we compare QSpec against individual quantization configurations (i.e., W4A4 and W4A16) and speculative decoding in terms of memory, computation, and generation. QSpec provides four key benefits: (1) Memory-Efficient. By sharing weights and overwriting KV caches, QSpec incurs memory costs on par with standalone high-precision activation quantization without any memory overhead caused by speculative decoding. (2) No Efficiency–Efficacy Trade-off. QSpec leverages speculative decoding to boost efficiency while preserving output quality, avoiding the compromises in conventional quantization. (3) Plug-and-Play Compatibility. QSpec requires only an acceptance policy and a cache-overwriting step, requiring no additional training or classifiers. Hence, it can be quickly integrated into existing quantization models. (4) High Acceptance Rate. Common weights and KV cache overwriting ensure consistent predictions, leading to a high rate of token acceptance.